OpenStudy (wow.):

WILL MEDAL!!! Which fraction is equivalent to -7/8?

OpenStudy (anonymous):

are there any options?

OpenStudy (wow.):

(A) 8/-7 (B) 7/-8 (C)-7/-8 (D)8/7

OpenStudy (bahrom7893):

-7/8=7/-8=-(7/8)

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Option A =>\[\frac{ 8 }{ -7 }=-\frac{ 8 }{ 7 }\]\[-\frac{ 7 }{ 8 }\neq-\frac{ 8 }{ 7 }\]Option B =>\[\frac{ 7 }{ -8 }=-\frac{ 7 }{ 8 }\]\[-\frac{ 7 }{ 8 }=-\frac{ 7 }{ 8 }\] Option C\[\frac{ -7 }{ -8 }=\frac{ 7 }{ 8 }\]\[\frac{ 7 }{ 8 }\neq- \frac{ 7 }{ 8 }\]Option D =>\[\frac{ 8 }{ 7 }\neq- \frac{ 7 }{ 8 }\]